# FOF-X: Towards Real-time Detailed Human Reconstruction from a Single Image

### v1

@@ Abstract
We introduce FOF-X for real-time reconstruction of detailed human geometry from a single image. Balancing real-time speed against high-quality results is a persistent challenge, mainly due to the high computational demands of existing 3D representations. To address this, we propose Fourier Occupancy Field (FOF), an efficient 3D representation by learning the Fourier series. The core of FOF is to factorize a 3D occupancy field into a 2D vector field, retaining topology and spatial relationships within the 3D domain while facilitating compatibility with 2D convolutional neural networks. Such a representation bridges the gap between 3D and 2D domains, enabling the integration of human parametric models as priors and enhancing the reconstruction robustness. Based on FOF, we design a new reconstruction framework, FOF-X, to avoid the performance degradation caused by texture and lighting. This enables our real-time reconstruction system to better handle the domain gap between training images and real images. Additionally, in FOF-X, we enhance the inter-conversion algorithms between FOF and mesh representations with a Laplacian constraint and an automaton-based discontinuity matcher, improving both quality and robustness. We validate the strengths of our approach on different datasets and real-captured data, where FOF-X achieves new state-of-the-art results. The code will be released for research purposes at https://cic.tju.edu.cn/faculty/likun/projects/FOFX/index.html.
@@ I Introduction
Reconstructing a 3D human from a single image has emerged as a popular task in computer vision and graphics, which can be widely used in various downstream applications, such as mixed reality and virtual try-on. However, real-time, high-fidelity monocular 3D human reconstruction remains challenging. The core of this challenge lies in the 3D representation, as it significantly influences the design and performance of reconstruction and generation approaches. Despite the promising results, existing methods [1, 2, 3, 4] typically suffer from high computational consumption and lack of robustness. Consequently, a good 3D geometry representation is essential for 3DTV, Holographic Telepresence systems, and other real-time applications, which must fulfill the requirements of accuracy, efficiency, compatibility, and generality [5].
@@
Implicit neural representations have emerged and been widely used in monocular human reconstruction [1, 11, 12, 13, 14, 3]. These methods treat 3D space as a continuous field, such as an occupancy or signed distance field, represented as $F(x,y,z):\mathbb{R}^{3}\to\mathbb{R}$. Instead of voxel grids, a neural network models the field, allowing results at any resolution. However, they infer values across numerous spatial grid points, which slows down inference and makes high-frame-rate or real-time reconstruction a significant challenge. Based on PIFu [1], Monoport [15] proposes an efficient sampling scheme to speed up inference, achieving 15 FPS only for mesh-free rendering. However, when generating a complete mesh, their method operates at less than 10 FPS, and the results are often worse than PIFu [1]. These limitations hinder its applicability with downstream applications.
@@
When directly applying FOF to a real-time system, we observed certain limitations in robustness. The reconstruction is degraded by texture and lighting effects. Additionally, efficiency and robustness issues in the original mesh-to-FOF conversion algorithm used in the conference version prevent the use of the SMPL prior in the real-time system, resulting in unsatisfactory performance on challenging poses. Furthermore, the original FOF-to-mesh algorithm comes with stair-step artifacts, further compromising the reconstruction quality. To address these challenges, we propose FOF-X, a novel reconstruction framework. With a single RGB image as input, FOF-X first translates it to dual-sided normal maps, which are unaffected by texture and lighting, as an internal representation of the human body. This allows the network to focus on geometric details, significantly improving the robustness and performance of the reconstruction. For the mesh-to-FOF conversion process, we designed an automaton-based discontinuity matcher to filter out illegal fragments, ensuring accurate FOF computations. This matcher can be parallelized on a GPU, enabling real-time conversion of parametric models. Additionally, we incorporate a Laplacian coordinate constraint in the FOF-to-mesh conversion process to eliminate artifacts caused by view bias, resulting in more accurate mesh reconstructions and improving the fidelity of the final output. As shown in Fig. 2, the newly designed mesh-to-FOF and FOF-to-mesh algorithms in FOF-X effectively address the robustness issues. Based on these updates, our final method, FOF-X, achieves new state-of-the-art results in both speed and accuracy, contributing a novel solution for real-time detailed human reconstruction from a single image. Experiments also show that FOF-X has better generalization for real-captured data and other objects.
@@
  • In FOF-X, we design parallelized inter-conversion algorithms between FOF and meshes with a Laplacian coordinate constraint and an automaton-based discontinuity matcher, further enhancing the robustness.
@@
An early version of this work has been published as a conference paper [20]. Based on our FOF representation, several works for generation[21] and reconstruction[22, 23] have been proposed. We have made significant extensions to our previous work in three aspects. Firstly, we design a new reconstruction framework greatly mitigating the performance degradation caused by texture and lighting effects. Secondly, we propose a robust mesh-to-FOF conversion algorithm with an automaton-based discontinuity matcher, enabling real-time execution and significantly improving the system’s robustness when facing challenging poses. Thirdly, we enhance the FOF-to-mesh algorithm with a Laplacian coordinate constraint for greater robustness and fidelity. Such a strategy effectively addresses artifacts caused by view direction bias without any loss of geometric details. Additionally, we conduct extensive experiments on FOF-X. The results from general object reconstruction and noise perturbation experiments demonstrate the strong generalization and robustness of FOF-X.
@@ II-A Surface-based Reconstruction
Parametric models, such as SMPL [16], SMPL-X [17], and STAR [18], are popular representations for naked 3D humans. These models are derived from large datasets of naked human shapes and use statistical methods to create analytical models, which can generate a human body mesh with just a few dozen parameters. Building on these models, many works [24, 25, 26] estimate 3D human shapes from an RGB image by predicting the parameters of the parametric model. However, these methods are limited in that they cannot reconstruct clothing or hair. More details can be found in Tian et al[27].
@@
Graph-based representations are naturally compatible with triangle meshes. Li et al[9] use graph neural networks to reconstruct topology-consistent geometries of clothed humans, but the recovered mesh tends to be smooth and lacks fine details. Representing 2D manifold surfaces of different 3D objects with a fixed structure is very challenging. Consequently, most methods heavily rely on the topology defined by SMPL and cannot produce detailed human geometries with complex topology. Habermann et al[29, 30] propose a deep learning approach for monocular dense human performance capture, but this method requires scanning the actor with a 3D scanner to create a template mesh. In a word, graph-based methods struggle to align image features accurately with the results, leading to unstable and over-smoothed outputs.
@@ II-B Volume-based Reconstruction
Recent advancements have seen the utilization of Neural Radiance Fields (NeRF) [34] in approaches like SHERF [35] and ELICIT [36]. NeRF applies volume rendering to implicit volume-based representations, enabling direct supervision of 3D scenarios with 2D images. Consequently, SHERF achieves the first generalizable Human NeRF from a single human image, while ELICIT employs a pre-trained CLIP model [37] for contextual understanding. However, NeRF-based approaches primarily focus on view synthesis and often lack detailed geometry generation from monocular images.
@@
In recent works [38, 39, 40, 41, 4], 3D objects are modeled using front and back depth maps, resembling a sandwich structure that encloses the inner space. However, depth maps only represent part of the geometric surface and cannot fully reconstruct the object. Therefore, these methods require volume-based post-processing, such as IF-Nets [42] and Poisson surface reconstruction[43], to complete the result further. Any-Shot GIN [41] extends the sandwich-like scheme to novel classes of objects by predicting the depth maps of two sides from a single image and performing shape completion with IF-Nets. ECON [4] follows a similar approach and extends it to 3D humans. It estimates the normal maps of two sides and converts them into depth maps using an optimization method. The parametric human model prior is also employed in the shape completion stage to “inpaint” the missing geometry. However, this optimization process is time-consuming, and its instability may lead to incorrect geometric estimates.
@@ II-D Leveraging Generative Models
Reconstructing a 3D human body from a single image is inherently challenging due to incomplete observations, making it difficult to regress geometric information accurately. In this case, generative models can offer plausible completion for invisible regions. SiTH [19] generates the back-view image with an image-conditioned diffusion module and then predicts dual-sided normal maps for subsequent reconstruction. TeCH [10] recovers human meshes with a SMPL-X initialized hybrid Deep Marching Tetrahedra (DMTet) representation and Score Distillation Sampling (SDS) Loss [44], achieving visually impressive results but with noisy surfaces. Human-LRM [45] first uses an LRM [46] model to reconstruct the human body as a NeRF from a single image. It then applies a diffusion model to refine the multi-view images rendered from the NeRF, followed by another LRM reconstruction based on the refined images to produce the final result. However, the heavy pipeline makes it inefficient for real-time use and does not provide a substantial improvement in reconstructed geometry quality. These generative model-based methods are typically time-consuming, and the reconstructed result does not accurately match the input image.
@@ III-A1 Vanilla FOF
where $\{a_{n}\},\{b_{n}\}\in\mathbb{R}$ are coefficients of basis functions $\left\{\cos(nx)\right\}$ and $\left\{\sin(nx)\right\}$, respectively. Note that by defining the occupancy function as Eq. (1), Eq. (2) holds for discontinuity at $z^{*}$ on the surface $S$ because $f(z^{*})=(f(z^{*}-)+f(z^{*}+))/2=0.5$, where $f(z^{*}-)$ and $f(z^{*}+)$ are the left hand limit and right hand limit of $f(z)$ at $z^{*}$. For compact representation, we approximate 1D occupancy signals by a subspace spanned by the first $12N+$ 1 basis functions:
@@
where $\bm{b}(z)=[1/2,\cos(z),\sin(z),\ldots,\cos(Nz),\sin(Nz)]^{\top}$ is the vector of the first $12N+$ 1 basis functions spanning the approximation subspace, and $\bm{c}=[a_{0},a_{1},b_{1},\ldots,a_{N},b_{N}]^{\top}$ is the coefficient vector which provides a more compact representation of the 1D occupancy function $f(z)$.
@@
where $\bm{C}(x,y)$ is the ($12N+$ 1)-dimensional Fourier coefficient vector for the 1D occupancy signal at $(x,y)$. In this way, we obtain the Fourier Occupancy Field $\bm{C}:[-1,1]^{2}\mapsto\mathbb{R}^{2N+1}$ for 3D occupancy field $F$.
@@ III-A2 FOF with Cosine Series
Because $f(z)$ is defined on $[-1,1]$ but not $[0,1]$, a shift on $z$ is needed, which contributes to the term $\frac{(z+1)}{2}$ in Eq. (5).
@@ III-B1 Reconstruction with Dual-sided Normal Maps
One important change in FOF-X is that we rely on dual-sided normal maps estimated from input RGB images rather than directly using the RGB images themselves to reconstruct human geometry. When testing with RGB input, we found a domain gap between the training images and real captured images, which significantly degraded the performance of our method. Although we try to mitigate this gap with a physical-based rendering engine (see Sec. IV-A), the reconstruction network is still sensitive to different textures and lighting conditions. To solve this, we design a new reconstruction framework with the dual-sided normal maps as the internal representation. Together with the FOF of the corresponding SMPL model, it is then fed to the network and used to reconstruct the human geometry. Since normal maps only keep the geometric information and are not affected by texture and lighting, the reconstruction network can focus on geometric information with better robustness and performance.
Thanks to the efficient dimensionality reduction with subspace approximation, the task of learning FOF from dual-sided normal maps can be essentially regarded as an image-to-image process. To this end, we exploit the image encoder building on the HR-Net [47] framework for its outstanding fitting capability in various vision tasks. We use the weak perspective camera model in our implementation, in which FOF (and thus the reconstructed geometry) is naturally aligned with the input RGB image. Note that FOF can also be used with the normalized device coordinate space (NDC space) so that various camera models, including the perspective camera, are compatible with the framework.
@@ III-C1 Mesh to FOF with Automaton-based Discontinuity Matcher
However, due to numerical errors and some non-watertight meshes, rasterization may produce unpaired $\mathcal{Z}$, i.e., a line may encounter successive discontinuities both going out of or into the mesh. And it can come to a going outside point first instead of a going inside point. To solve this, we propose an automaton-based discontinuity matcher to enhance the robustness. Unlike what we do in the conference version [20], we collect those discontinuities and normal vectors for each line first and process them with an automaton before working out the integral. Let’s consider the discontinuities corresponding to one pixel. We sort these points by $z$ first and then represent them with a sequence of 00s and $1$s, where 00 stands for the normal vector facing front and $1$ stands for the normal vector facing backwards. We add a stop letter \$currency-dollar\\$\$ to the sequence and then process it with the automaton shown in Fig. 5, which effectively handles irregular patterns (initial $1$s and repeating 00s or $1$s). After being filtered by the automaton, the produced sequence can then be integrated with Eq. (8).

### v2

@@ Abstract
We introduce FOF-X for real-time reconstruction of detailed human geometry from a single image. Balancing real-time speed against high-quality results is a persistent challenge, mainly due to the high computational demands of existing 3D representations. To address this, we propose Fourier Occupancy Field (FOF), an efficient 3D representation by learning the Fourier series. The core of FOF is to factorize a 3D occupancy field into a 2D vector field, retaining topology and spatial relationships within the 3D domain while facilitating compatibility with 2D convolutional neural networks. Such a representation bridges the gap between 3D and 2D domains, enabling the integration of human parametric models as priors and enhancing the reconstruction robustness. Based on FOF, we design a new reconstruction framework, FOF-X, to avoid the performance degradation caused by texture and lighting. This enables our real-time reconstruction system to better handle the domain gap between training images and real images. Additionally, in FOF-X, we enhance the inter-conversion algorithms between FOF and mesh representations with a Laplacian constraint and an automaton-based discontinuity matcher, improving both quality and robustness. We validate the strengths of our approach on different datasets and real-captured data, where FOF-X achieves new state-of-the-art results. The code has already been released for research purposes at https://cic.tju.edu.cn/faculty/likun/projects/FOFX/index.html.
@@ I Introduction
Reconstructing a 3D human from a single image has emerged as a popular task in computer vision and graphics, which can be widely used in various downstream applications, such as mixed reality and virtual try-on. However, real-time, high-fidelity monocular 3D human reconstruction remains challenging. The core of this challenge lies in the 3D representation, as it significantly influences the design and performance of reconstruction and generation approaches. Despite some promising results, existing methods [1, 2, 3, 4] typically suffer from high computational consumption and lack of robustness. Consequently, a good 3D geometry representation is essential for 3DTV, Holographic Telepresence systems, and other real-time applications, which must fulfill the requirements of accuracy, efficiency, and compatibility [5].
@@
Implicit neural representations have emerged and been widely used in monocular human reconstruction [1, 11, 12, 13, 3]. These methods treat 3D space as a continuous field, such as an occupancy or signed distance field, represented as $F(x,y,z):\mathbb{R}^{3}\to\mathbb{R}$. Instead of voxel grids, a neural network models the field, allowing results at any resolution. However, these methods infer values across numerous spatial grid points, which slows down inference and makes high-frame-rate or real-time reconstruction a significant challenge. Based on PIFu [1], Monoport [14] proposes an efficient sampling scheme to speed up inference, achieving 15 FPS only for mesh-free rendering. However, when generating a complete mesh, their method operates at less than 10 FPS, and the results are often worse than PIFu. These limitations hinder Monoport’s applicability with downstream applications.
@@
When applying FOF directly to a real-time system, we observed certain limitations in robustness. The reconstruction is degraded by texture and lighting effects. Additionally, efficiency and robustness issues in the original mesh-to-FOF conversion algorithm used in the conference version prevent the use of the SMPL prior in the real-time system, resulting in unsatisfactory performance on challenging poses. Furthermore, the original FOF-to-mesh algorithm comes with stair-step artifacts, further compromising the reconstruction quality. To address these challenges, we propose FOF-X, a novel reconstruction framework. With a single RGB image as input, FOF-X first translates it to dual-sided normal maps, which are unaffected by texture and lighting, as an internal representation of the human body. This strategy allows the network to focus on geometric details, significantly improving the robustness and performance of the reconstruction. To ensure the robustness and fidelity of FOF computations, we design an automaton-based discontinuity matcher to filter out invalid fragments. This matcher can be parallelized on a GPU, enabling real-time conversion of parametric models. Additionally, we incorporate a Laplacian coordinate constraint in the FOF-to-mesh conversion process to eliminate artifacts caused by view bias, resulting in more accurate mesh reconstructions and improving the fidelity of the final output. As shown in Fig. 2, the newly designed mesh-to-FOF and FOF-to-mesh algorithms in FOF-X effectively address the robustness issues. Based on these updates, our final method, FOF-X, achieves new state-of-the-art results in both speed and accuracy, contributing a novel solution for real-time detailed human reconstruction from a single image. Experiments also show that FOF-X has better generalization for real-captured data.
@@
  • In FOF-X, we design parallelized inter-conversion algorithms between FOF and meshes with a Laplacian coordinate constraint and an automaton-based discontinuity matcher, further enhancing the robustness and reconstruction quality.
@@
An early version of our method, including only FOF, was published as a conference paper [19]. Based on our FOF representation, several works for generation[20] and reconstruction[21, 22, 23] have been proposed. In this paper, we substantially extend the original version by developing FOF-X. Firstly, we design a new reconstruction framework that greatly mitigates the performance degradation caused by texture and lighting effects. Secondly, we propose a robust mesh-to-FOF conversion algorithm with an automaton-based discontinuity matcher, enabling real-time execution and significantly improving the system’s robustness when facing challenging poses. Thirdly, we propose a FOF-to-mesh algorithm with a Laplacian coordinate constraint for greater robustness and fidelity. Such a strategy effectively addresses artifacts caused by view-direction bias without any loss of geometric details. Fourth, we employ cosine series as the subspace approximation basis, retaining the Fourier-like approximation capacity while eliminating Gibbs phenomena through even-periodic boundary extension. Additionally, we evaluate FOF-X through comprehensive experiments including: testing on more datasets, comparisons with various state-of-the-art methods, and more detailed ablation studies, all confirming the method’s effectiveness, efficiency and robustness.
FOF-X offers three key advantages: (1) FOF-X is sampling-scalable along all three spatial axes, allowing for resolution adaptation at inference without retraining, which supports diverse deployment scenarios with varying speed, quality, and memory constraints; (2) FOF-X represents 3D objects as multi-channel images, requiring only a simple tensor multiplication to reconstruct the 3D occupancy field; (3) FOF-X can be seamlessly inter-converted to and from the mesh representation, making it compatible with traditional graphics pipelines. Additionally, as a compact and computationally efficient 3D representation, FOF-X is suitable for real-time applications like holographic transportation.
@@ II-A Surface-based Reconstruction
Parametric models, such as SMPL [15], SMPL-X [16], and STAR [17], are popular representations for naked 3D humans. These models are derived from large datasets of naked human shapes and use statistical methods to create analytical models, which can generate a human body mesh with just a few dozen parameters. Building on these models, many works [24, 25] estimate 3D human shapes from an RGB image by predicting the parameters of the parametric model. However, these methods are limited in that they cannot reconstruct clothing or hair.
@@
Graph-based representations are naturally compatible with triangle meshes. Li et al[9] use graph neural networks to reconstruct topology-consistent geometries of clothed humans, but the recovered mesh tends to be smooth and lacks fine details. Representing 2D manifold surfaces of different 3D objects with a fixed structure is very challenging. Consequently, most methods heavily rely on the topology defined by SMPL and cannot produce detailed human geometries with complex topology. Habermann et al[27] propose a deep learning approach for monocular dense human performance capture, but this method requires scanning the actor with a 3D scanner to create a template mesh. In a word, graph-based methods struggle to align image features accurately with the results, leading to unstable and over-smoothed outputs.
@@ II-B Volume-based Reconstruction
Recent advancements have seen the utilization of Neural Radiance Fields (NeRF) [31] in various reconstruction approaches, and subsequent methods [32, 33, 34] have been further adapted for human reconstruction. NeRF applies volume rendering to implicit volume-based representations, enabling direct supervision of 3D scenarios with 2D images. Consequently, SHERF [33] achieves the first generalizable Human NeRF from a single human image, while ELICIT [34] employs a pre-trained CLIP model [35] for contextual understanding. However, NeRF-based approaches primarily focus on view synthesis and often lack detailed geometry generation from monocular images.
@@
In recent works [36, 37, 4], 3D objects are modeled using front and back depth maps, resembling a sandwich structure that encloses the inner space. However, depth maps only represent part of the geometric surface and cannot fully reconstruct the object. Therefore, these methods require volume-based post-processing, such as IF-Nets [38] and Poisson surface reconstruction[39], to complete the result further. Any-Shot GIN [37] extends the sandwich-like scheme to novel classes of objects by predicting the depth maps of two sides from a single image and performing shape completion with IF-Nets. ECON [4] follows a similar approach and extends it to 3D humans. It estimates the normal maps of two sides and converts them into depth maps using an optimization method. The parametric human model prior is also employed in the shape completion stage to “inpaint” the missing geometry. However, this optimization process is time-consuming, and its instability may lead to incorrect geometric estimates.
@@ II-D Leveraging Generative Models
Reconstructing a 3D human body from a single image is inherently challenging due to incomplete observations, making it difficult to regress geometric information accurately. In this case, generative models can offer plausible completion for invisible regions. SiTH [18] generates the back-view image with an image-conditioned diffusion module and then predicts dual-sided normal maps for subsequent reconstruction. TeCH [10] recovers human meshes with a SMPL-X initialized hybrid Deep Marching Tetrahedra (DMTet) representation and Score Distillation Sampling (SDS) Loss [40], achieving visually impressive results but with noisy surfaces. Human-LRM [41] first uses an LRM [42] model to reconstruct the human body as a NeRF from a single image. It then applies a diffusion model to refine the multi-view images rendered from the NeRF, followed by another LRM reconstruction based on the refined images to produce the final result, but does not provide a substantial improvement in reconstructed geometry quality. PSHuman [43] first uses a multi-view image diffusion model to generate six views of global full-body images and local face images. Then, using the generated normal and color maps, SMPL-X is deformed and remeshed by differentiable rasterization. However, the heavy pipeline makes it inefficient for real-time use. These generative model-based methods are typically time-consuming, and the reconstructed result does not accurately match the input image.
@@ III-A1 Vanilla FOF
[…]
@@
[…]
@@
[…]
@@ III-A2 FOF with Cosine Series
Because $f(z)$ is defined on [1,1][-1,1][ - 1 , 1 ] but not [0,1][0,1][ 0 , 1 ], a shift on $z$ is needed, which contributes to the term $\frac{(z+1)}{2}$ in Eq. (5).
@@ III-B Image-to-Image Reconstruction Network
In FOF-X, we introduce a novel reconstruction network with dual-sided normal map to mitigate lighting interference in real-world capture scenarios, coupled with integrated SMPL priors that enhance pose robustness under challenging articulations.
@@ III-B1 Reconstruction with Dual-sided Normal Maps
One important change in FOF-X is that we rely on dual-sided normal maps estimated from input RGB images rather than directly using the RGB images themselves to reconstruct human geometry. When testing with RGB input, we found a domain gap between the training images and real captured images, which significantly degraded the performance of our method. Although we try to mitigate this gap with a physics-based rendering engine (see Sec. IV-A), the reconstruction network is still sensitive to different textures and lighting conditions. To solve this, we design a new reconstruction framework with the dual-sided normal maps as the internal representation. Together with the FOF of the corresponding SMPL model, it is then fed to the network and used to reconstruct the human geometry. Since normal maps only keep the geometric information and are not affected by texture and lighting, the reconstruction network can focus on geometric information with better robustness and performance.
Thanks to the efficient dimensionality reduction with subspace approximation, the task of learning FOF from dual-sided normal maps can be essentially regarded as an image-to-image process. To this end, we exploit the image encoder built on the HR-Net [44] framework for its outstanding fitting capability in various vision tasks. We use the weak perspective camera model in our implementation, in which FOF (and thus the reconstructed geometry) is naturally aligned with the input RGB image. Note that FOF can also be used with the normalized device coordinate space (NDC space) so that various camera models, including the perspective camera, are compatible with the framework.
@@ III-C1 Mesh to FOF with Automaton-based Discontinuity Matcher
Figure 6: Example of Discontinuity Point Matching via Automaton.
However, due to numerical errors and some non-watertight meshes, rasterization may produce unpaired $\mathcal{Z}$, i.e., a line may encounter successive discontinuities both going out of or into the mesh. And it can come to a going outside point first instead of a going inside point. To solve this, we propose an automaton-based discontinuity matcher to enhance the robustness. Unlike what we do in the conference version [19], we collect those discontinuities and normal vectors for each line first and process them with an automaton before working out the integral. Let’s consider the discontinuities corresponding to one pixel. We sort these points by $z$ first and then represent them with a sequence of 0s and 111s, where 0 stands for the normal vector facing front and 111 stands for the normal vector facing backwards. We add a stop letter \$\\$\$ to the sequence and then process it with the automaton shown in Fig. 5, which effectively handles irregular patterns (initial 111s and repeating 0s or 111s). After being filtered by the automaton, the produced sequence can then be integrated with Eq. (8). For watertight meshes, both our automaton and the conference version produce identical results, correctly converting meshes to FOF. For non-watertight meshes (e.g., surfaces with open boundaries or self-intersections, partly missing, or with non-manifold connectivity), the conference version method fails to compute FOF accurately, leading to floating artifacts (Fig. 2). Our automaton-based discontinuity matcher robustly handles irregular patterns arising from non-watertight meshes (including holes, self-intersections, and duplicate faces), effectively preventing reconstruction artifacts caused by such topological imperfections. Fig. 6 shows an example of our automaton processing a non-watertight mesh.